# Rainbow PO: A Unified Framework for Combining Improvements in Preference Optimization

### v1

@@ 3.1 Component Descriptions
$=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}^{*}}{\mathbb{E}}\log\sigma\left(\frac{\beta}{|y^{w}|}\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\gamma}\left(y_{w}\mid x\right)}-\frac{\beta}{|y^{l}|}\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\gamma}\left(y_{l}\mid x\right)}\right).$
@@
This transformation motivates us to further propose a new mechanism which we call as mixing reference policy. If taking $\pi_{\text{sft}}$ as the reference policy is too conservative (not strong enough), and taking $\pi_{\gamma}$ policy can help improve the performance but totally neglects the original SFT model implicit preference, can we benefit from a mixing of these two policies? The answer is YES. Consider a exponential mixing (or a linear combination of the loglikehood) with $\alpha\in[0,1]$, defined as:
$\pi_{\alpha}(y\mid x)\propto\pi^{\alpha}_{\text{ref}}(y\mid x)\cdot\pi^{1-\alpha}_{\gamma}(y\mid x).$ (10)
@@
If we use $\pi_{\alpha}$ as the reference policy, we can write $\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{\alpha})$ as a pratical form:
@@
$-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)^{1/|y^{w}|}}{\pi_{\theta}\left(y_{l}\mid x\right)^{1/|y^{l}|}}-\alpha\beta\log\frac{\pi_{\text{ref }}\left(y_{w}\mid x\right)^{1/|y^{w}|}}{\pi_{\text{ref }}\left(y_{l}\mid x\right)^{1/|y^{l}|}}-(1-\alpha)\gamma\right).$ (11)
Notice that $\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{\alpha})=\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{0})=\mathcal{L}_{\text{SimPO}}(\pi_{\theta})$, thus SimPO is an instance of mixing policy by taking $\alpha=0$; $\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{1})=\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta})$, thus $\alpha=1$ corresponds to DPO applied with length normalization. Because of convexity of $-\log(\sigma(\cdot))$, the mixing policy objective can also be understood as a lower bound of the linear combination of the LN-DPO loss and SimPO loss:
$\mathcal{\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi^{\alpha})}\leq\alpha\cdot\mathcal{\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{\text{ref}})}+(1-\alpha)\mathcal{\mathcal{L}_{\text{SimPO}}(\pi_{\theta};\pi_{\text{ref}})}.$ (12)
@@
According to our experiment results, we indeed find that there exists $\alpha\in(0,1)$ that performs better than both sides, see Figure 1(b). Recent work (Liu et al., 2024b) analyze the role of reference model, and argue that stronger reference model could benefit DPO; our finding is consistent, as we further explicitly design a choice of better reference model for better performance.
@@
Since the proposal of DPO, there is controversy on the exact equivalence of DPO and RLHF. RSO (Liu et al., 2023), further pointed out that the data should be generated from the optimal policy if treating DPO objective as maximum likelihood estimation. Thus RSO adopts a statistical rejection sampling for sampling preference dataset generated by the optimal policy to mitigate this distribution difference in DPO.
@@
To address the intrinstic different variance schedules of reward model for different prompts, and stablize the process for formulating the preference dataset, we also adopt a modified version of RSO by computing the percentile reward (or the ranking reward) in the whole generation set instead of utilizing the true reward, which we found that can stabilize the generation and yield better results when further applied with DPO, as in Algorithm 1. Similar to in RSO (Liu et al., 2023), we search the best temperature hyper-parameter for RSO through the downstream task performance as the validation metric, which we detail in Figure 1(c). We then use the empirically best performed temperature constant $\tau$ to formulate the prefer-
@@
Existing work also considered the contextual difference: some preference pairs might be of higher uncertainty or have more dispersion. In this paper, we adopt the idea of Mallows-DPO by introducing a contextual scaling factor $\phi(x)$ on the likelihood difference, e.g., for DPO:
$-\log\sigma\left(\phi(x)\left[\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right]\right),$ (13)
which is motivated by the Mallows ranking model and the form of Mallows-DPO (Chen et al., 2024a). In Mallows-DPO, $\phi(x)$ corresponds to a normalized predictive entropy of the preference pair $(x,y^{w},y^{l})$:
$\phi(x)=-\log\left(\frac{\sum_{i=1}^{N-1}\left[H_{\pi_{\mathrm{ref}}}\left(Y_{i+1}\mid Y_{i}=y_{i}^{w}\right)+H_{\pi_{\mathrm{ref}}}\left(Y_{i+1}\mid Y_{i}=y_{i}^{l}\right)\right]}{2\log n}\right).$ (14)
@@
SFT loss is straightforward by adding extra SFT loss term on the winning answer, or a reference answer (for regularization, which appears for reference-free methods like CPO and ORPO); however, according to our experiments, we find that adding SFT loss could largely degrade the performance.
@@ 3.2 Unified Formulation: Rainbow
$\mathcal{L}_{\textsc{RainbowPO}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right):=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}^{*}}{\mathbb{E}}f\left[\phi(x)\left(\frac{\beta}{|y^{w}|^{\eta}}\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\alpha}\left(y_{w}\mid x\right)}-\frac{\beta}{|y^{l}|^{\eta}}\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\alpha}\left(y_{l}\mid x\right)}\right)\right],$ (15)
in which $\eta\in\{0,1\}$, $\pi_{\alpha}$ is as defined in Equation 10. The preference dataset $\mathcal{D}^{*}$ can be $\mathcal{D}_{\text{RS}}$, if we have access to the reward model’s true value, which means that the preference dataset is formulated by by rejection sampling from the original dataset’s prompts as mentioned in Algorithm 1. If the reward model is black-box oracle, namely we cannot access the true reward value, we will always utilize the usual formulation way of preference dataset $\mathcal{D}$, with the details in the Experiment Section.
Like all other XPO methods, to achieve a final satisfactory performance, RainbowPO can introduce an extensive amount of hyper-parameter search for the best performing $f$, $\alpha$, $\beta$, $\gamma$ and whether $\eta=1$. For efficient hyper-parameter search, we conducted a greedy search method with the help of our framework and decomposition of effective elements: we search for the best hyper-parameters for those that affects the performance in the most when we gradually add designs to the preference optimization methods. For example, when adding length normalization to the methods, we only search for the best hyper-parameter for the regularization parameter $\beta$, and will fix the learning rate and all the training args, which prevents the parameter searching space from exploding.
@@ 4 Experiments
We choose Llama3-8B-Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct as our model base to fine tune, mainly because that aligning this widely adopted and flagship instruct model is of great interest to the whole community and meets the standard as a representative setup for alignment. It can also help mitigate the uncertainty from probably not perfectly supervised fine-tuned models.
For evaluation metric, we use widely adopted benchmark Alpaca Eval2, which is composed of 805 questions and evaluate the instruction following capability of the model. The win rate is by default annotated by GPT4 through LLM-as-a-judge, and the resulting win rate has a 68.5% consistency according to official AlpacaEval website. To cross validate the effectiveness of the model and mitigate possible bias of GPT4, we also adopt Llama3-70B instruct as the judge, which is reported to have a 67.5% win rate consistency to humans.
For formulating the preference dataset $\mathcal{D}$, we follow the standard RLHF pipeline by directly generating answers from the model (which is thus an on-policy dataset, but the algorithm is still offline) and get AI feedbacks as in SimPO (Meng et al., 2024): we generate 5 answers from Llama3-8B-Instruct for each prompt in UltraFeedback (Cui et al., 2023), rank them with scores evaluated by ArmoRM (Wang et al., 2024a), and choose the best/worst one as winning/losing answer to form the preference pairs. For training, we adopted the popular library Transformer Reinforcement Learning (TRL [^3] https://huggingface.co/docs/trl/index), which already implemented most aforementioned xPOs algorithms and make everything under the same backend and easy to reproduce. If not specified, we train the model with 3 training epochs, which typically yields better performance for each xPOs according to our replication.
@@ 4.1 Effectiveness of Different Components
We first study the effectiveness of adding individual components. We use $+$ to denote that only the component(s) after is added to DPO baseline as in Table 2. From the results, we could notice that some components may not provide firm improvement over the baseline not matter being added individually or combined. For example, for home advantage term, we tune different values under the best performed $\beta$ for DPO, and also always witness a degradation in the performance, see Figure 1(a). For link function, we examine the square loss in IPO and also cannot see performance gain over the DPO baseline. Other components (LN, Mixing reference policy, CS) indeed help improve the metric even added individually. Compared to SimPO, using mixing reference policy yields also better results as in Figure 1(b). The average win rate gain is reported in the last column.
Models AlpacaEval (GPT4) AlpacaEval (Llama3-70B)
@@
Table 2: Model performance results on each component. (based on training-epochs = 3)
@@
Figure 1: Investigation on the dynamics of changing home advantage, reference policy mixing and different temperature in RSO.
@@
Models AlpacaEval (GPT4) AlpacaEval (Llama3-70B)
@@
We also conduct an ablation study of our proposed RainbowPO algorithm. All components of our proposed in our algorithm is useful, as in Table 4, for which we use $\oplus$ to denote that the methods are based on composition of the method on previous line and new elements in Table 4. We notice that adding length normalization is indeed important and of the most critical importance among the components for RainbowPO.
Models AlpacaEval (GPT4) AlpacaEval (Llama3-70B)
@@
Table 4: Ablation study of the newly proposed elements in RainbowPOwithout the use of a trained reward model.
@@ 4.2 Comparison with Baseline Methods
Models AlpacaEval (GPT4) AlpacaEval (Llama3-70B)
@@
Models AlpacaEval (GPT4) AlpacaEval (Llama3-70B)
@@ 4.3 Limitations and Future Work
In this paper, we focus our evaluation on models trained with LLama3-8B Instruct as the base model. Exploring other models of varying sizes, such as Gemma (Team et al., 2024) or Mistral (Jiang et al., 2023a), could possibly enhance the generalizability of our findings. It will also be beneficial if we could repeat the pipelines and compare the algorithms’ performance on other LLM evaluation metrics, like arena-hard or MT-bench, though MT-bench is known to be less tinguishable for RLHF algorithms. Other directions include benchmarking the effectiveness of alignment algorithms on improving other capabilities of LLM other than instruction following, like reasoning Xiong et al. (2024b). However, due to constraints in computing resources and time, we defer this investigation to future work. Nevertheless, we believe that our work provides a unified and comprehensive framework for helping to find the best preference optimization algorithms, and further pushing the boundary of offline RLHF for LLMs.
@@
We were not able to explore other aspects of existing DPO variants in detail, and there might be still promising candidates in further improving the preference of RainbowPO. Some methods that propose to update the reference policy dynamically: sDPO (Kim et al., 2024), TR-DPO (Gorbatovski et al., 2024). Additionally, we also recognize the recent literature in pursuing online methods, such as online DPO (Guo et al., 2024) or iterative DPO (Yuan et al., 2024; Xiong et al., 2024a), which provide valuable insights on possibly further improving the downstream task performance: we will pursue them in future research. Other extensions beyond RLHF include, Nash Learning from human feedback (Munos et al., 2023), and self-play preference optimization (Chen et al., 2024b).
@@
We also made some interesting observations in the paper, which we fail to find proper mathematical explanations and may boost further research. For example, the RainbowPO objective could benefit much more than SimPO objective when increasing the training epochs, but the mathematical reasons for such phenomenons are still unknown. In addition, we found some mathematically orthogonal components are actually not empirically independent, for example, RSO can improve DPO, but can not be readily combined with other components like length normalization. It is also interesting to see the some combination of components reach effects “$1+1>2$”; it will be interesting to understand the deeper underlying reasons and could potentially lead to better algorithms.

### v2

@@ 3.1 Component Descriptions
$\mathcal{L}_{\mathrm{SimPO}}\left(\pi_{\theta};\gamma\right)\equiv-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\log\sigma\left(\frac{\beta}{|y_{w}|}\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\gamma}\left(y_{w}\mid x\right)}-\frac{\beta}{|y_{l}|}\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\gamma}\left(y_{l}\mid x\right)}\right).$ (12)
@@
This transformation motivates us to further propose a new mechanism which we call as mixing (reference) policy. If taking $\pi_{\text{sft}}$ as the reference policy is too conservative (not strong enough), and taking $\pi_{\gamma}$ policy can help improve the performance but totally neglects the original SFT model implicit preference information, can we benefit from a mixing of these two policies? The answer is YES. Consider a exponential mixing of likelihood of reference policy $\pi_{\gamma}$ and $\pi_{\text{ref}}$ (we use $\pi_{\alpha,\gamma}$ to denote the relevance of the resulting reference policy to both $\alpha$ and $\gamma$) with $\alpha\in[0,1]$, defined as:
$\pi_{\alpha,\gamma}(y\mid x)\propto\pi^{\alpha}_{\text{ref}}(y\mid x)\cdot\pi^{1-\alpha}_{\gamma}(y\mid x).$ (13)
@@
Then if we use $\pi_{\alpha,\gamma}$ as the reference policy in (7), we yield $\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{\alpha,\gamma})$ with a practical form:
@@
$-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)^{1/|y_{w}|}}{\pi_{\theta}\left(y_{l}\mid x\right)^{1/|y_{l}|}}-\alpha\beta\log\frac{\pi_{\text{ref }}\left(y_{w}\mid x\right)^{1/|y_{w}|}}{\pi_{\text{ref }}\left(y_{l}\mid x\right)^{1/|y_{l}|}}-(1-\alpha)\gamma\right).$ (14)
Notice that $\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{\alpha,\gamma})=\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{0,\gamma})=\mathcal{L}_{\text{SimPO}}(\pi_{\theta};\gamma)$, thus SimPO is an instance of mixing policy by taking $\alpha=0$; $\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta};\pi_{1,\gamma})=\mathcal{L}_{\text{LN-DPO}}(\pi_{\theta})$, thus $\alpha=1$ corresponds to DPO applied with length normalization as in Equation 7. For finding a good $\pi_{\alpha,\gamma}$, we first take $\alpha=0$ and tune the best $\gamma$, which is similar to SimPO; we then afterwards tune $\alpha$ using obtained $\gamma$.
@@
According to our experiment results, we indeed find that there exists $\alpha\in(0,1)$ that performs better than both sides (i.e. $\alpha=0$ or $\alpha=1$), see Figure 1(b). Recent work (Liu et al., 2024b) analyze the role of reference model, and argue that stronger reference model could benefit DPO; our finding is consistent, as we further explicitly design a choice of better reference model for better performance.
@@
Since the proposal of DPO, there is controversy on the exact equivalence of DPO and RLHF. RSO (Liu et al., 2023), further pointed out that the data should be generated from the optimal policy if treating DPO objective as maximum likelihood estimation. Thus RSO adopts a statistical rejection sampling for sampling preference dataset $\mathcal{D}$ generated by the optimal policy to mitigate this distribution difference in DPO.
@@
To address the intrinstic different variance schedules of reward model for different prompts, and stablize the process for formulating the preference dataset, we also adopt a modified version of RSO by computing the percentile reward (or the ranking reward) in the whole generation set instead of utilizing the true reward, which we found that can stabilize the generation and yield better results when further applied with DPO, as in Algorithm 1.
Similar to in RSO (Liu et al., 2023), we search the best temperature hyper-parameter for RSO through the downstream task performance as the validation metric, which we detail in Figure 1(c). We then use the empirically best performed temperature constant $\tau$ to formulate the prefer-
@@
Existing work also considered the contextual difference: some preference pairs might be of higher uncertainty or have more dispersion. In this paper, we adopt the idea of MallowsPO in Chen et al. (2024a) by introducing a contextual scaling factor $\phi(x)$ on the likelihood difference, which yields:
$\mathcal{L}_{\text{MallowsPO}}(\pi_{\theta};\pi_{\text{ref}})=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\log\sigma\left(\phi(x)\left[\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right]\right),$ (15)
if adding this factor to DPO objective. The scaling factor is motivated by the Mallows ranking model which has a natural carrier of dispersion index. In MallowsPO, $\phi(x)$ corresponds to a normalized predictive entropy of the preference pair $(x,y_{w},y_{l})$ ($y_{w,i}$ and $y_{l,i}$ denote the $i^{\text{th}}$ component of $y_{w}$ and $y_{l}$):
$\phi(x)=-\log\left(\frac{\sum_{i=1}^{N-1}\left[H_{\pi_{\mathrm{ref}}}\left(Y_{i+1}\mid Y_{i}=y_{w,i}\right)+H_{\pi_{\mathrm{ref}}}\left(Y_{i+1}\mid Y_{i}=y_{l,i}\right)\right]}{2\log n}\right).\vspace{-5pt}$ (16)
@@
SFT loss is straightforward by adding extra SFT loss term on the winning answer, or a reference answer (for regularization, which appears for reference-free methods like CPO and ORPO). For example, if adding SFT for DPO with preferred/winning answers, we yield:
$\mathcal{L}_{\text{SFT}}\left(\pi_{\theta};\lambda\right):=-\mathbb{E}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\text{ref }}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\text{ref }}\left(y_{l}\mid x\right)}\right)+\lambda\log(\pi_{\theta}\left(y_{w}\mid x\right))\right],$ (17)
However, we find that adding SFT loss could largely degrade the performance, as in Section 4.
@@ 3.2 Unified Formulation: Rainbow
$\mathcal{L}_{\textsc{RainbowPO}}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right):=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}^{*}}{\mathbb{E}}f\left[\phi(x)\left(\frac{\beta}{|y_{w}|^{\eta}}\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\alpha,\gamma}\left(y_{w}\mid x\right)}-\frac{\beta}{|y_{l}|^{\eta}}\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\alpha,\gamma}\left(y_{l}\mid x\right)}\right)\right],$ (18)
in which $\eta\in\{0,1\}$, and $\pi_{\alpha,\gamma}$ is as defined in Equation (13). The preference dataset $\mathcal{D}^{*}$ can be $\mathcal{D}_{\text{RS}}$, which means that the preference dataset is formulated by by rejection sampling from the original dataset’s prompts as mentioned in Algorithm 1, if we have access to a reward model’s true value. If the reward model is black-box oracle, namely we cannot access the true reward value, we will always utilize the usual formulation way of preference dataset $\mathcal{D}$, detailed in the experiment section.
Hyper-parameters. Like all other xPOs, to achieve the best performance, RainbowPO can introduce an extensive amount of hyper-parameter search for the best performing $f$, $\alpha$, $\beta$, $\gamma$ and whether $\eta=1$. For efficient hyper-parameter search, we conducted a greedy search method with the help of our framework and decomposition of effective elements: we search for the best hyper-parameters for those that affects the performance in the most when we gradually add designs to the preference optimization methods. For example, when adding length normalization to the methods, we only search for the best hyper-parameter for the regularization parameter $\beta$, and will fix the learning rate and all the training args, which prevents the parameter searching space from exploding.
@@ 4 Experiments
We choose Llama3-8B-Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. as our model base to fine tune, mainly because that aligning this widely adopted and flagship instruct model is of great interest to the whole community and meets the standard as a representative setup for alignment. It can also help mitigate the uncertainty from probably not perfectly supervised fine-tuned models.
For evaluation metric, we use widely adopted benchmark AlpacaEval2, which is composed of 805 questions and evaluate the instruction following capability of the model. AlpacaEval2 evaluates models with the win rates (WR) of model generations against the reference/base answers generated by GPT4-turbo. The comparisons are by default annotated by a GPT4-turbo and the resulting WR has a 68.5% consistency with human evaluation, according to official AlpacaEval2 website. Length Controlled (LC) Win Rate (WR) is a debiased version of the WR that control for the length of the outputs and increase the WR’s correlation to Chat Arena, [^3] https://lmarena.ai/. while significantly decreasing the length gameability of the annotator. To cross validate the effectiveness of the model and mitigate possible bias of GPT4, we also adopt Llama3-70B instruct as the judge, which is reported to have a 67.5% win rate consistency to humans (close to the performance of GPT4). We include more detailed background information of AlpacaEval2 in Appendix E.1.
For formulating the preference dataset $\mathcal{D}$, we follow the standard RLHF pipeline by directly generating answers from the model (which is thus an on-policy dataset, but the algorithm is still offline) and get AI feedbacks as in SimPO (Meng et al., 2024): we generate 5 answers from Llama3-8B-Instruct for each prompt in UltraFeedback (Cui et al., 2023), rank them with scores evaluated by ArmoRM (Wang et al., 2024a), and choose the best/worst one as winning/losing answer to form the preference pairs. For training, we adopted the popular library Transformer Reinforcement Learning (TRL), [^4] https://huggingface.co/docs/trl/index. which already implemented most aforementioned xPOs algorithms and make everything under the same backend and easy to reproduce. If not specified, we train the model with 3 training epochs, which typically yields better performance for each xPOs according to our replication.
@@ 4.1 Effectiveness of Different Components
We first study the effectiveness of adding individual components. We use $+$ to denote that only the component(s) after is added to DPO baseline as in Table 2. From the results, we could notice that some components may not provide firm improvement over the baseline, no matter being added individually or combined. For example, for home advantage, we tune different values under the best performed $\beta$ for DPO, and also always witness a degradation in performances, see Figure 1(a). For link function, we examine the square loss in IPO and did not see performance gain over the DPO baseline. Other components (LN, Mixing reference policy, CS) indeed help improve the metric even added individually. Compared to SimPO, using mixing reference policy yields also better results as in Figure 1(b). The average WR gain is reported in the last column.
Models AlpacaEval2 (GPT4) AlpacaEval2 (Llama3-70B)
@@
Table 2: Model performance results on each component after model training for 3 epochs.
@@
Figure 1: Dynamics of changing home adv., reference policy mixing and different temp. in RSO.
@@
Models AlpacaEval2 (GPT4) AlpacaEval2 (Llama3-70B)
@@
We also conduct an ablation study of our proposed RainbowPO algorithm. All components of our proposed in our algorithm is useful, as in Table 4, for which we use $\oplus$ to denote that the methods are based on composition of the method on previous line and new elements in Table 4. We notice that adding length normalization is indeed important and of the most critical importance among the components for RainbowPO. We also include ablations on training epochs in Appendix E.3, which showcases that 3 epochs yield the sweet spot.
Models AlpacaEval2 (GPT4) AlpacaEval2 (Llama3-70B)
@@
Table 4: Ablation study of the newly proposed elements in RainbowPO.
@@ 4.2 Comparison with Baseline Methods
Models AlpacaEval2 (GPT4) AlpacaEval2 (Llama3-70B)
@@
Models AlpacaEval2 (GPT4) AlpacaEval2 (Llama3-70B)
@@ 4.3 Limitations and Future Work
In this paper, we focus our evaluation on models trained with LLama3-8B Instruct as the base model. Exploring other models of varying sizes, such as Gemma (Team et al., 2024) or Mistral (Jiang et al., 2023a), could possibly enhance the generalizability of our findings. It will also be beneficial if we could repeat the pipelines and compare the algorithms’ performance on other LLM evaluation metrics, like arena-hard or MT-bench, though MT-bench is known to be less tinguishable for RLHF algorithms. Other directions include benchmarking the effectiveness of alignment algorithms on improving other capabilities of LLM other than instruction following, like reasoning (Xiong et al., 2024b) or COT (Choi et al., 2024). However, due to constraints in computing resources and time, we defer this investigation to future work. Nevertheless, we believe that our work provides a unified and comprehensive framework for helping to find the best preference optimization algorithms, and further pushing the boundary of offline RLHF for LLMs.
@@
We were not able to explore other aspects of existing DPO variants in detail, and there might be still promising candidates in further improving the preference of RainbowPO. Some methods that propose to update the reference policy dynamically: sDPO (Kim et al., 2024), TR-DPO (Gorbatovski et al., 2024); or learning from noisy preferences Chowdhury et al. (2024). Additionally, we also recognize the recent literature in pursuing online methods, such as online DPO (Guo et al., 2024) or iterative DPO (Yuan et al., 2024; Xiong et al., 2024a), which provide valuable insights on possibly further improving the downstream task performance: we will pursue them in future research. Other extensions beyond RLHF include, Nash Learning from human feedback (Munos et al., 2023), and self-play preference optimization (Chen et al., 2024b).
@@
We also made some interesting observations in the paper, which we fail to find proper mathematical explanations and may boost further research. For example, the RainbowPO objective could benefit much more than SimPO when increasing the training epochs, but reasons for such phenomenons are still unknown. In addition, we found some mathematically orthogonal components are actually not empirically independent, for example, RSO can improve DPO, but can not be readily combined with other components like length normalization. It is also interesting to see the some combination of components reach effects “$1+1>2$”; it will be interesting to understand the deeper underlying reasons and could potentially lead to better algorithms.